OpenStudy (anonymous):

For a school drama performance, student tickets cost $5 each and adult tickets cost $10 each. The sellers collected $3,570 from 512 tickets sold. If c is the number of student tickets sold, which equation can be used to find the amount of tickets sold of each type? A. 10(512 – c) + 5c = 512 B. 10(512 – c) + 5c = 3,570 C. 5(512 – c) + 10c = 3,570 D. 5(512 – c) + 10c 512

OpenStudy (anonymous):

5c+10a=3570 c+a=512 ... a=512-c so 5c + 10(512-c)=3570

OpenStudy (anonymous):

A B C or D?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thank you!! (:

OpenStudy (anonymous):

No prob!